# Visual CoT: Unleashing Chain-of-Thought Reasoning in Multi-Modal Language Models

### v1

@@ 3.3 Framework
To train the MLLM with visual CoT data, we add a CoT prompt (“Please provide the bounding box coordinate of the region that can help you answer the question better.”) to the question, asking the model to identify the most informative region of the image. VisCoT then determines this region and generates its bounding box (bbox). Using this bbox and the original image, a visual sampler extracts the localized image $X_{1}$ containing detailed information. The same vision encoder and projector are used to extract visual tokens $H_{1}$. The MLLM then integrates data from both the original and localized images {$H_{0}$, $H_{1}$} to provide more precise and comprehensive answers. For data without visual CoT annotations, this procedure is omitted as indicated by the dashed box in Fig. 3. Here, the MLLM directly answers based on the input image alone. Our VisCoT model is thus adaptable to data in both annotated and non-annotated formats simultaneously.
@@
Visual Sampler. The visual sampler’s role is to accurately select the relevant region from the queried image. We first calculate the center point $[x_{0},y_{0}]$, the half-width $w_{half}$, and the half-height $h_{half}$ of the bounding box predicted by the MLLM. To capture more contextual information and meet the square receptive field requirement of the CLIP model, $\max\{\max\{w_{half},h_{half}\},\text{res}_{half}\}$ is chosen as the sample size $s$. $\text{res}_{half}$ is the half input size of the vision encoder. Consequently, the visual sampler crops the region $[x_{0}-s,y_{0}-s,x_{0}+s,y_{0}+s]$ for further processing. During inference, if the calculated cropped box extends beyond the image boundaries, the center point is adjusted towards the center of the image to ensure the box remains within the image frame. This adjustment is important for improving the overall performance, as it can mitigate the impact of any detection inaccuracies.
Inference. During inference, VisCoT offers two methods for generating answers: with or without the visual CoT (Chain of Thought) process. To engage the CoT feature, the visual CoT prompt should be added to the question. Differing from the training stage, we further include the prompt “Please answer the question based on the original image and local detail image. [Question]” after presenting the localized image $X_{1}$, where “[Question]” is the original question. This approach encourages the MLLM to focus more effectively on both the original and localized images while revisiting the initial question, leading to improved performance. Besides, this additional prompt helps the model to better integrate and consider the information from both image sources, resulting in more accurate and comprehensive answers.
3.4 Model Training
@@
MLLM Res. Flickr30k GQA Open images VSR Birds-200-2011
LLaVA-1.5-7B [32] 3362 0.581 0.534 0.412 0.572 0.530 0.444
LLaVA-1.5-13B [32] 3362 0.620 0.571 0.413 0.590 0.573 0.468
SPHINX-13B [30] 2242 0.607 0.584 0.467 0.613 0.505 0.407
VisCoT-7B 2242 0.671 0.616 0.833 0.682 0.556 0.547
VisCoT-7B 3362 0.668 0.631 0.822 0.614 0.559 0.582
Table 2: Performance on the Visual CoT benchmark. Datasets highlighted in grey indicate that their respective training splits were not used in our model’s training phase. Res indicates input image resolution.
VisCoT is trained in two stages. In the first stage, consistent with LLaVA-1.5, we freeze the weights of the vision encoder and LLM and utilize image-text caption data for training. In the second stage, all weights within our framework are trainable. We train the model on a reorganized Vision-Language dataset. The training data is a composite of three sources: the second stage data from LLaVA, data from Shikra’s [6] second stage, and our visual CoT data. For more detailed information on the data composition, please refer to the appendix. The inclusion of data from Shikra, which features various datasets with positional annotations, such as RefCOCO [18] for REC, visual gemone [21] for grounding caption. These datasets can enhance VisCoT’s ability to accurately identify and understand locations within images. This enhancement is crucial for tasks requiring precise spatial awareness.
@@ 4 Experiments
Firstly, we provide an overview of the training details of VisCoT. Subsequently, in the evaluation phase, we begin by accessing VisCoT on traditional multimodal and grounding benchmarks (refer to Sec. 4.1). Additionally, we conduct further experiments to analyze the impact of essential components within VisCoT through an ablation study in Sec. 4.2. Finally, we showcase the capabilities of VisCoT in engaging complex multimodal conversations in Sec. 4.3.
Training Details. Following the setup described by Vicuna [10], our model undergoes a two-stage training process. In the first stage, we pre-train the model for 1 epoch using a learning rate of 2e-3 and a batch size of 128. For the second stage, we fine-tune the model for 1 epoch on our visual CoT dataset, employing a learning rate of 2e-5 and a batch size of 128. The Adam optimizer with zero weight decay and a cosine learning rate scheduler are utilized. To conserve GPU memory during fine-tuning, we employ FSDP (Full Shard Data Parallel) with ZeRO3-style. All models are trained using 32 $\times$ A100s. In the case of training the setting with a 7B LLM and a resolution of 224, the first/second pre-training stage completes within 1/16 hours.
Res. 3362 2242 2242 3362 3362
DocVQA 21.6 14.4 39.0 29.4 49.3
TextVQA 58.2 55.5 62.9 60.2 66.9
ChartQA 17.7 14.2 19.2 17.5 22.8
Table 3: Performance on VQA benchmarks.
[…]
BLIP-2 [25] Vicuna-13B 2242 41.0 42.5 85.3 1293.8
InstructBLIP [11] Vicuna-7B 2242 49.2 50.1 36.0
InstructBLIP [11] Vicuna-13B 2242 49.5 50.7 78.9 1212.8
Shikra [6] Vicuna-13B 2242 58.8
IDEFICS-9B [24] LLaMA-7B 2242 44.2 38.4 25.9 48.2
IDEFICS-80B [24] LLaMA-65B 2242 68.9 45.2 30.9 54.5
Qwen-VL [3] Qwen-7B 4482 67.1 59.3 63.8 38.2
Qwen-VL-Chat [3] Qwen-7B 4482 68.2 57.5 61.5 1487.5 360.7 60.6
LLaVA1.5 [33] Vicuna-7B 3362 66.8 62.0 58.2 85.9 1510.7 64.3
LLaVA1.5 [33] Vicuna-13B 3362 71.6 63.3 61.3 85.9 1531.3 295.4 67.7
SPHINX [3] LLaMA-13B 2242 69.3 62.6 51.6 80.7 1476.1 310.0 66.9
VisCoT Vicuna-7B 2242 69.2 62.9 55.8 86.1 1437.6 285.7 66.6
VisCoT Vicuna-13B 2242 71.6 64.2 57.8 85.6 1480.0 255.4 66.9
VisCoT Vicuna-7B 3362 68.3 62.0 61.0 86.5 1514.4 275.0 67.3
VisCoT Vicuna-13B 3362 73.6 63.3 62.3 83.3 1535.7 331.8 67.5
[…]
@@ 4.1 Performance evaluation
In this section, we present a comprehensive evaluation of VisCoT across various multi-modal tasks to thoroughly assess our model’s visual understanding ability. Tab. 2 highlights the enhancements through the visual CoT benchmark. In Tab. 4 and Tab. 5, we showcase the baseline performance of our model, where it directly answers questions without employing the visual CoT process.
Visual CoT Benchmark. In Tab. 2, we test our model and LLaVA-1.5 on the proposed visual CoT benchmark as detailed in Sec. 3.2. To demonstrate the impact of the chain-of-thought process, we also include the ablation study that removes this reasoning process and directly generates the response in a standard, direct manner. Notably, we find that our pipeline shows significant improvement in the doc/text-related tasks and high-resolution image processing. This is evident even in cases where the training splits from some datasets were not been utilized for the model training. For instance, SROIE [16] is a dataset consisting of scanned receipt images and we need to extract the key information from them, such as the company name and the total price. Our model achieves 8$\times$ performance compared to the standard pipeline without the chain of thought. Furthermore, the visual CoT pipeline also shows superior results in other benchmark tasks, showing its efficacy in enhancing the model’s comprehensive visual and textual interpretation abilities.
Multi-modal Large Language Models Benchmarks. In Tab. 4, we evaluate our model on recently proposed MLLM benchmarks such as MME [12], POPE [29], MMbench [35], ScienceQA [36], TextVQA [50], GQA [17]. Our model still achieves comparative results across all benchmarks. This performance indicates that the visual CoT data we proposed not only enhances visual comprehension in CoT-specific scenarios but also boosts the model’s overall visual understanding in standard inference setups. As demonstrated in Tab. 3, the implementation of visual CoT enables our model to achieve superior performance even with a lower resolution and a reduced number of visual tokens. This finding highlights the efficiency and effectiveness of the visual CoT approach in enhancing model accuracy.
@@ 4.2 Ablation study
Flickr30k GQA Open images VSR Birds-200-2011
Baseline 0.671 0.616 0.833 0.682 0.556 0.547
w/o CoT 0.610 0.600 0.656 0.634 0.534 0.433
GT BBox 0.692 0.796 0.896 0.792 0.577 0.757
Random 0.627 0.477 0.763 0.585 0.683 0.453
Center 0.653 0.547 0.803 0.657 0.609 0.487
Table 6: Ablation study on the different BBox selection strategies. ‘w/o CoT’ indicates a standard, non-CoT-based inference process. ‘GT BBox’ denotes we replace the predicted bboxes with the annotated ground truth bboxes. ‘Random’ and ‘Center’ refer to using random and center bboxes instead of model predictions.
@@
0.461 0.356 0.671 0.710 0.556 0.547
Table 7: Ablation study on the visual sampler design.
Figure 5: Ablation study on the visual CoT prompt design.
Visual CoT Prompt Design. As illustrated in Fig. 5, we conducted ablation experiments to optimize the visual CoT prompt design. Unlike previous ablations, the benchmarks in this table are from official sources. From Type1 to Type2, we found that repeating the original question after presenting the localized image enhances accuracy. Further, by introducing an additional prompt to direct the MLLM’s focus to both images, we observed a subsequent improvement in accuracy. This indicates the effectiveness of prompt design in guiding the MLLM’s attention and improving its performance in visual CoT tasks.
@@ 4.3 Visualization
This section displays VisCoT’s qualitative performance through Fig. 6, highlighting its visual CoT ability to identify critical regions in images that aid in answering questions and synthesizing the combined contexts of both original and zoomed-in images. We also provide comparative results with different configurations: VisCoT (GT BBox), and VisCoT (w/o CoT), which are defined in Tab. 6. We find that the accuracy of detection and depth of understanding directly contribute to the quality of the generated answers.
@@ 5 Conclusion
In this paper, we introduced VisCoT, a pioneering approach that enhances multi-modal large language models (MLLMs) with visual Chain-of-Thought reasoning. This methodology addresses critical gaps in MLLMs, particularly in interpretability and processing dynamic, focused visual inputs. Our proposed visual CoT dataset offers 373k annotated question-answer pairs for detailed visual analysis. Our novel multi-turn processing pipeline allows MLLMs to dynamically focus and interpret visual data, mirroring human cognition. Meanwhile, VisCoT provides more interpretable reasoning stages. The introduction of the visual CoT benchmark is a step forward in evaluating MLLMs’ ability to concentrate on specific image areas. The extensive experiments validate the framework’s effectiveness. Our work serves as an encouraging starting point for further exploration and development in the field of visual CoT.
Supplementary material of Visual CoT: Unleashing Chain-of-Thought Reasoning in the Multi-Modal Language Model
Appendix 0.A Prompt design
0.A.1 Generating the dataset for TextCaps
You are an AI visual assistant, and you are seeing a single image. What you see is provided with several sentences and Ocr_tokens, describing the same image you are looking at. Ocr_tokens indicates the text in the image. Answer all questions as you are seeing the image. Design a conversation between you and a person asking about this photo. The answers should be in a tone that a visual AI assistant is seeing the image and answering the question. Ask THREE diverse questions and give corresponding answers. Again, do not ask about uncertain details. Do not just makeup questions and answers based on Ocr tokens. Your response should include questions asking about the textual information of the image, the object types, counting the objects, object actions, object locations, relative positions between objects, etc. Please only ask questions that have definite answers: One can see the content in the image that the question asks about and can answer confidently; One can determine confidently from the image that it is not in the image. Do not ask any questions that cannot be answered confidently. One can not see the Ocr_tokens, so the question must not mention ‘Ocr’ Craft Questions Around Ocr_tokens: Create questions that directly pertain to these identified words or phrases. Ensure that the question is structured in a way that the answer MUST be a word or phrase directly from the Ocr_tokens. Your answer cannot contain words outside of Ocr_tokens. The answers must be within three words. Please follow the provided format: Question: [question] Answer: [answer] Here is the context you need to process: Image description: %s Ocr_tokens: %s
0.A.2 Generating the dataset for Flickr30k
0.A.3 Evaluation for the visual CoT benchmark using the ChatGPT
Question: %s Standard answer: %s Model’s answer: %s
Appendix 0.B Training data details
@@
We listed all training data in Table 8. We removed the images from the training set that are the same as those in the testing or validation set to prevent potential data leakage. Our training data includes three parts, and they are from LLaVA-1.5, a subset of Shikra, and our proposed visual CoT dataset separately.
@@
Visual CoT dataset 373K TextVQA, TextCaps, DocVQA, Birds-200-2011
@@
Table 8: The overview of our training dataset.
Appendix 0.C Detection performance of the visual CoT bboxes
@@
MLLM Res. Flickr30k GQA Open images VSR Birds-200-2011
VisCoT-7B 2242 49.6 42.0 57.6 69.6 67.0 37.8
VisCoT-7B 3362 51.3 49.5 59.3 54.0 47.1 38.2
VisCoT-13B 2242 48.7 38.7 53.8 53.5 49.6 34.4
Appendix 0.D Limitation

### v2

@@ 4 Enhancing MLLMs with Chain-of-Thought Capabilities
VisCoT Pipeline. To train the MLLM baseline with visual CoT data, we add a CoT prompt (“Please provide the bounding box coordinate of the region that can help you answer the question better.”) to the question, asking the model to identify the most informative region of the image. VisCoT then determines this region and generates its bounding box. During the training phase, we utilize the ground truth bounding box to extract visual information rather than a predicted one in the following steps. With the original image $X_{0}$ and the bbox, a visual sampler extracts the localized image $X_{1}$ containing detailed information. The same vision encoder and projector are then used to extract visual tokens $H_{1}$. The MLLM then integrates visual tokens from both the original and localized images {$H_{0}$, $H_{1}$} to provide more precise and comprehensive answers. For data without visual CoT annotations, this procedure is omitted as indicated by the dashed box in Fig. 3. Here, the MLLM directly answers based on the input image alone. Our VisCoT baseline is thus adaptable to data in both annotated and non-annotated formats simultaneously.
Figure 3: VisCoT first extracts visual tokens from an image and pinpoints the key region relevant to the question. Then, it processes the localized visual information. Finally, the MLLM integrates the information from the overall and localized images to construct a comprehensive and accurate answer.
@@
Visual Sampler. Given the original image and the predicted bbox, the visual sampler’s role is to accurately select the relevant region that considers the visual encoder requirement and bbox corner cases. We first calculate the center point $[x_{0},y_{0}]$, half-width $w_{half}$, and half-height $h_{half}$ of the bounding box predicted by VisCoT. To capture more context and meet the square receptive field requirement of the CLIP model, $\max\{\max\{w_{half},h_{half}\},\text{res}_{half}\}$ is chosen as the sample size $s$. $\text{res}_{half}$ is the half input size of the vision encoder. Consequently, the visual sampler crops the region $[x_{0}-s,y_{0}-s,x_{0}+s,y_{0}+s]$ for further processing. During inference, if the calculated cropped box extends beyond the image boundaries, the center point is adjusted towards the center of the image to ensure the box remains within the image frame. This adjustment is important for improving the overall performance, as it can mitigate the impact of any detection inaccuracies.
Inference. VisCoT offers two options to generate answers: with or without the visual CoT process. If the CoT feature is not needed, users can simply provide the MLLM with the image and question. To engage the CoT feature, users can append the additional visual CoT prompt after the question.
Model Training VisCoT baseline is trained in two stages. In the first stage, consistent with LLaVA-1.5, we freeze the weights of the vision encoder and LLM, and utilize image-text caption data for training. In the second stage, all weights are trainable. For more details, see Appendix B.
Table 3: Performance on the Visual CoT benchmark. Datasets highlighted in grey indicate their training splits were not used in our model’s training phase. Res indicates input image resolution.
@@
MLLM Res. Flickr30k Visual7W GQA Open images VSR Birds-200-2011
LLaVA-1.5-7B [33] 3362 0.581 0.575 0.534 0.412 0.572 0.530 0.454
LLaVA-1.5-13B [33] 3362 0.620 0.580 0.571 0.413 0.590 0.573 0.478
SPHINX-13B [31] 2242 0.607 0.558 0.584 0.467 0.613 0.505 0.419
VisCoT-7B 2242 0.671 0.580 0.616 0.833 0.682 0.556 0.550
VisCoT-7B 3362 0.668 0.558 0.631 0.822 0.614 0.559 0.580
@@ 5 Experiments
Firstly, we provide an overview of the construction and evaluation of the CoT benchmark. Subsequently, in the evaluation phase, we begin by accessing VisCoT on the proposed benchmark (refer to Sec. 5.2). Additionally, we conduct further experiments to analyze the impact of essential components within VisCoT through an ablation study in Sec. 5.3. Finally, we showcase the capabilities of VisCoT in engaging complex multimodal conversations in Sec. 5.4. The training details and detection performance of the visual CoT bboxes can be found in Appendix B & C.
5.1 Visual CoT Benchmark
In this section, we provide an overview of our visual CoT benchmark, which primarily focuses on scenarios where the MLLM needs to concentrate on specific regions within a complete image. We utilize 12 source datasets, as shown in Fig. 1, and when an official training/evaluation split exists, we adopt it. In cases where such a split does not exist, we randomly divide the dataset. Additionally, we incorporate the test split of SROIE, DUDE, and Visual7W to evaluate the model’s zero-shot visual CoT capabilities. Following the methodology of previous MLLM studies [28, 40], we employ ChatGPT [45] and ask it to assign a numerical score between 0 and 1, where a higher score indicates better prediction accuracy. For detailed information on the prompt used for ChatGPT-based evaluation, please refer to Appendix E.4.
@@ 5.2 Performance Evaluation
In this section, we comprehensively evaluate VisCoT across various multi-modal tasks to thoroughly assess our model’s visual understanding ability. Tab. 3 highlights the enhancements through the visual CoT benchmark. We also showcase the baseline performance of our model on other benchmarks in Appendix D, where it directly answers questions without employing the visual CoT process.
In Tab. 3, we test our model and LLaVA-1.5 on the proposed visual CoT benchmark as detailed in Sec. 5.1. To demonstrate the impact of the chain-of-thought process, we also include the ablation study that removes this reasoning process and directly generates the response in a standard, direct manner. Notably, our pipeline shows significant improvement in the doc/text-related tasks and high-resolution image processing, even when the training splits from corresponding datasets are not utilized for the model training. For instance, SROIE [17] is a dataset that involves extracting key information from scanned receipts, such as the company name and the total price. Our model achieves 8$\times$ performance compared to the standard pipeline without a chain-of-thought process. Furthermore, the visual CoT pipeline also shows superior results in other benchmark tasks, showing its efficacy in enhancing the model’s comprehensive visual and textual interpretation abilities.
@@ 5.3 Ablation Study
Table 4: Ablation study on the different BBox selection strategies. ‘w/o CoT’ indicates a standard, non-CoT-based inference process. ‘GT BBox’ uses annotated ground truth bboxes. ‘Random’ and ‘Center’ refer to using random and center bboxes instead of model predictions.
@@
Flickr30k Visual7W GQA Open images VSR Birds-200-2011
Baseline 0.671 0.580 0.616 0.833 0.682 0.556 0.550
w/o CoT 0.610 0.554 0.600 0.656 0.634 0.534 0.443
GT BBox 0.692 0.699 0.796 0.896 0.792 0.577 0.752
Random 0.627 0.458 0.477 0.763 0.585 0.683 0.453
Center 0.653 0.529 0.547 0.803 0.657 0.609 0.490
@@
Table 5: Ablation study on the visual sampler design.
@@
0.461 0.356 0.626 0.710 0.556 0.550
@@ 5.4 Visualization
This section displays VisCoT’s qualitative performance through Fig. 4, highlighting its visual CoT ability to identify critical regions in images that aid in answering questions and synthesizing the combined contexts of both original and zoomed-in images. We also provide comparative results with different configurations: VisCoT (GT BBox), and VisCoT (w/o CoT). The accuracy of detection and depth of understanding directly contribute to the quality of the generated answers.
@@ 6 Conclusion
In this paper, we introduced VisCoT, a pioneering approach that enhances multi-modal large language models with visual chain-of-thought reasoning. This methodology addresses critical gaps in MLLMs, particularly in interpretability and processing dynamic visual inputs. Our visual CoT dataset offers 438k annotated question-answer pairs for detailed visual analysis. Our novel multi-turn processing pipeline allows MLLMs to dynamically focus and interpret visual data, mirroring human cognition. VisCoT provides more interpretable reasoning stages, and the visual CoT benchmark advances the evaluation of MLLMs’ focus on specific image areas. Extensive experiments validate the framework’s effectiveness, offering a promising starting point for further exploration in visual CoT.
Appendix A Overview
Our supplementary includes the following sections:
  • Section B: Framework details. Details for model design, implementation and training data.
  • Section C: Detection performance of the visual CoT bboxes. Details for detection performance for the intermediate visual CoT bounding boxes.
  • Section D: More experiment results. Additional performance evaluation and performance analysis.
  • Section E: Prompt design. Prompt for generating the visual CoT dataset and evaluating the performance.
  • Section F: Limitations. Discussion of limitations of our work.
  • Section G: Potential negative societal impacts. Discussion of potential negative societal impacts of our work.
  • Section F: More visualization. More Visualization of our dataset and demos.
  • Section I: Disclaimer. Disclaimer for the visual CoT dataset and the related model.
Following NeurIPS Dataset and Benchmark track guidelines, we have shared the following artifacts:
Artifcat Link License
Code Repository https://github.com/deepcs233/Visual-CoT Apache-2.0 license
Data https://huggingface.co/datasets/deepcs233/Visual-CoT CC BY 4.0
Model Weights https://huggingface.co/collections/deepcs233/viscot-65fe883e2a0cdd3c59fc5d63 Apache-2.0 license
The authors are committed to ensuring its regular upkeep and updates.
Appendix B Framework details
B.1 Model details
We choose the pre-trained ViT-L/14 of CLIP [50] as the vision encoder and Vicuna-7/13B [10] as our LLM, which has better instruction following capabilities in language tasks compared to LLaMA [56]. Consider an input original image, we take the vision encoder to obtain the visual feature. Similar to LLaVA [34, 33], we use a simple linear layer to project the image features into the word embedding space to obtain the visual tokens $H_{0}$ which share the same dimensionality of the LLM.
B.2 Implementation details
Following the setup described by Vicuna [10], our model undergoes a two-stage training process. In the first stage, we pre-train the model for 1 epoch using a learning rate of 2e-3 and a batch size of 128. For the second stage, we fine-tune the model for 1 epoch on our visual CoT dataset, employing a learning rate of 2e-5 and a batch size of 128. The Adam optimizer with zero weight decay and a cosine learning rate scheduler are utilized. To conserve GPU memory during fine-tuning, we employ FSDP (Full Shard Data Parallel) with ZeRO3-style. All models are trained using 32 $\times$ A100s. In the case of training the setting with a 7B LLM and a resolution of 224, the first/second pre-training stage completes within 1/16 hours.
B.3 Training data details
@@
We train the model on a reorganized Vision-Language dataset. The training data is a composite of three sources: the second stage data from LLaVA, data from Shikra’s [6] second stage, and our visual CoT data. The inclusion of data from Shikra, which features various datasets with positional annotations, such as RefCOCO [19] for REC, visual gemone [22] for grounding caption. These datasets can enhance VisCoT’s ability to accurately identify and understand locations within images. This enhancement is crucial for tasks requiring precise spatial awareness. We listed all training data in Table 6. We removed the images from the training set that are the same as those in the testing or validation set to prevent potential data leakage. Our training data includes three parts, and they are from LLaVA-1.5, a subset of Shikra, and our proposed visual CoT dataset separately.
Table 6: The overview of our training dataset.
@@
Visual CoT dataset 376K TextVQA, TextCaps, DocVQA, Birds-200-2011
@@
Appendix C Detection performance of the visual CoT bboxes
@@
MLLM Res. Flickr30k Visual7W GQA Open images VSR Birds-200-2011
VisCoT-7B 2242 49.6 31.1 42.0 57.6 69.6 67.0 37.2
VisCoT-7B 3362 51.3 29.4 49.5 59.3 54.0 47.1 37.6
[…]
BLIP-2 [26] Vicuna-13B 2242 41.0 42.5 85.3 1293.8
InstructBLIP [11] Vicuna-7B 2242 49.2 50.1 36.0 23.7
InstructBLIP [11] Vicuna-13B 2242 49.5 50.7 78.9 1212.8
Shikra [6] Vicuna-13B 2242 58.8
IDEFICS-9B [25] LLaMA-7B 2242 44.2 38.4 25.9 48.2 25.2
IDEFICS-80B [25] LLaMA-65B 2242 68.9 45.2 30.9 54.5 38.1
Qwen-VL [3] Qwen-7B 4482 67.1 59.3 63.8 38.2 7.4
Qwen-VL-Chat [3] Qwen-7B 4482 68.2 57.5 61.5 1487.5 360.7 60.6 56.7
LLaVA1.5 [34] Vicuna-7B 3362 66.8 62.0 58.2 85.9 1510.7 64.3 58.3
LLaVA1.5 [34] Vicuna-13B 3362 71.6 63.3 61.3 85.9 1531.3 295.4 67.7 63.6
SPHINX [3] LLaMA-13B 2242 69.3 62.6 51.6 80.7 1476.1 310.0 66.9 56.2
VisCoT Vicuna-7B 2242 68.2 63.1 55.4 86.0 1453.6 308.3 67.9 59.7
VisCoT Vicuna-13B 2242 71.6 64.2 57.8 85.6 1480.0 255.4 66.9 60.5
VisCoT Vicuna-7B 3362 68.3 62.0 61.0 86.5 1514.4 275.0 67.3 60.1
VisCoT Vicuna-13B 3362 73.6 63.3 62.3 83.3 1535.7 331.8 67.4 61.6
Table 10: Performance on VQA benchmarks.
Res. 3362 2242 2242 3362 3362
DocVQA 21.6 14.4 39.0 29.4 49.3
TextVQA 58.2 55.5 62.9 60.2 66.9
ChartQA 17.7 14.2 19.2 17.5 22.8
Appendix D More experiment results
D.1 Performance evaluation
In Tab. 8 and Tab. 9, we showcase the baseline performance of our model, where it directly answers questions without employing the visual CoT process. Multi-modal Large Language Models Benchmarks. In Tab. 8, we evaluate our model on recently proposed MLLM benchmarks such as MME [13], POPE [30], MMbench [36], ScienceQA [37], TextVQA [53], GQA [18]. Our model still achieves comparative results across all benchmarks. This performance indicates that the visual CoT data we proposed not only enhances visual comprehension in CoT-specific scenarios but also boosts the model’s overall visual understanding in standard inference setups. As demonstrated in Tab. 10, the implementation of visual CoT enables our model to achieve superior performance even with a lower resolution and a reduced number of visual tokens. This finding highlights the efficiency and effectiveness of the visual CoT approach in enhancing model accuracy.
D.2 Performance analysis
Tab. 4 shows that our baseline with visual CoT performs better than the model without CoT. We further investigate whether different bounding box sizes affect performance improvement. In Fig. 6, we divide each evaluation dataset into five equal parts based on their relative bounding box sizes. We observe that the visual CoT usually achieve greater improvement when the corresponding bounding box is relative smaller.
Figure 6: Visualization of performance improvement across different bounding box relative sizes for different source datasets. We find that visual CoT shows a larger improvement in cases where the queried object is relatively small. Red bars represent evaluation data samples where the model with CoT outperforms the model without CoT. Green bars indicate the opposite. The y-axis represents different ranges of relative sizes of bboxes $R$. For example, the 20-40% range indicates that the bboxes in this range occupy the relatively small 20-40% quantile within the entire dataset. For clarity, samples where both models achieve the same scores are omitted.
Appendix E Prompt design
E.1 Generating the dataset for TextCaps
You are an AI visual assistant, and you are seeing a single image. What you see is provided with several sentences and Ocr_tokens, describing the same image you are looking at. Ocr_tokens indicates the text in the image. Answer all questions as you are seeing the image. Design a conversation between you and a person asking about this photo. The answers should be in a tone that a visual AI assistant is seeing the image and answering the question. Ask THREE diverse questions and give corresponding answers. Again, do not ask about uncertain details. Do not just makeup questions and answers based on Ocr tokens. Your response should include questions asking about the textual information of the image, the object types, counting the objects, object actions, object locations, relative positions between objects, etc. Please only ask questions that have definite answers: One can see the content in the image that the question asks about and can answer confidently; One can determine confidently from the image that it is not in the image. Do not ask any questions that cannot be answered confidently. One can not see the Ocr_tokens, so the question must not mention ‘Ocr’ Craft Questions Around Ocr_tokens: Create questions that directly pertain to these identified words or phrases. Ensure that the question is structured in a way that the answer MUST be a word or phrase directly from the Ocr_tokens. Your answer cannot contain words outside of Ocr_tokens. The answers must be within three words.
Please follow the provided format: Question: [question] Answer: [answer]
Here is the context you need to process: Image description: { } Ocr_tokens: { }
E.2 Generating the dataset for Flickr30k
E.3 Generating the dataset with detailed reasoning steps for GQA
You are an AI visual assistant, and you are seeing a single image. I will provide a question-answer pair along with the corresponding reasoning steps. The question and answer are based on an image. You need to generate the pure reasoning text in a step-by-step format, with each step clearly numbered (1. 2. 3. … etc). The reasoning text should help solve the question and reach the final answer without including or hinting at the answer itself. The reasoning text must not include any ID numbers.
Here is the data you need to process:
Question: What appliance is to the right of the cabinet?
Answer: The appliance is a microwave.
Reasoning steps: [{"operation": "select", "dependencies": [], "argument": "cabinet (3588933)"}, {"operation": "relate", "dependencies": [0], "argument": "appliance,to the right of,s (1564001)"}, {"operation": "query", "dependencies": [1], "argument": "name"}]
$\cdots$
E.4 Evaluation for the visual CoT benchmark using the ChatGPT
Question: { } Standard answer: { } Model’s answer: { }
Appendix F Limitations